# Learning Disentangled Identifiers for Action-Customized Text-to-Image Generation

### v1

@@ Abstract
This study focuses on a novel task in text-to-image (T2I) generation, namely action customization. The objective of this task is to learn the co-existing action from limited data and generalize it to unseen humans or even animals. Experimental results show that existing subject-driven customization methods fail to learn the representative characteristics of actions and struggle in decoupling actions from context features, including appearance. To overcome the preference for low-level features and the entanglement of high-level features, we propose an inversion-based method Action-Disentangled Identifier (ADI) to learn action-specific identifiers from the exemplar images. ADI first expands the semantic conditioning space by introducing layer-wise identifier tokens, thereby increasing the representational richness while distributing the inversion across different features. Then, to block the inversion of action-agnostic features, ADI extracts the gradient invariance from the constructed sample triples and masks the updates of irrelevant channels. To comprehensively evaluate the task, we present an ActionBench that includes a variety of actions, each accompanied by meticulously selected samples. Both quantitative and qualitative results show that our ADI outperforms existing baselines in action-customized T2I generation.
@@
*Work done during internship at Alibaba ${}^{\dagger}$orresponding author.
@@ 1 Introduction
To avoid the appearance leakage while accurately modeling the target action, we propose Action-Disentangled Identifier (ADI) to learn the optimal action-specific identifiers. Firstly, we expand the semantic conditioning space by applying layer-wise identifier tokens. Since existing works have analyzed that different layers have varying degrees of control over low-level and high-level features [29], such an expansion increases the accommodation of various features, making it easier to invert action-related features. Furthermore, we would like to decouple the action-agnostic features from the learning of action identifiers. To achieve this, we discover invariant mechanisms in the data that are difficult to vary across examples. Specifically, given an exemplar image with the specific action, another same-action image can be randomly sampled from the training data, forming a context-different tuple. Meanwhile, leveraging mature subject-driven customization techniques, an image that shares similar context can be quickly synthesized to form an action-different tuple. To decouple the highly-coupled features, we disentangle action-agnostic features at the gradient level, and construct two context gradient masks by comparing the difference on the gradients over the input tuples. By overwriting the merged gradient mask to the gradient of the anchor image, the update of action-agnostic channels on the identifiers is discarded.
@@
  • We propose a novel action customization task, which requires to learn the desired action from limited data for future generation. While existing customization focus on reprinting appearances, we highlight this under-studied but important problem.
@@ 2.1 Text-to-Image (T2I) Generation
Generating high-quality and diverse images from textual conditions has received considerable attention from both the research community and the general public. The previous dominant generative adversarial networks (GANs) [20, 30, 28, 37], consisting of a generator and a discriminator, suffer from unstable optimization and less diverse generations due to the adversarial training [2]. And variational autoencoders (VAEs) [8], which apply a probabilistic encoder-decoder architecture, are also prone to posterior collapse and over-smoothed generations [36]. Text-conditional auto-regressive models [18, 4, 3, 33] have shown more impressive results, but require time-consuming iterative processes to achieve high-quality image sampling. More recently, diffusion models have emerged as a promising alternative, achieving impressive results with open-vocabulary text descriptions through their natural fitting to inductive biases of image data [26, 19, 15, 22]. GLIDE [15] introduces text conditions into the diffusion process through the use of an unclassified guide. DALL-E 2 [19] employs a diffusion prior module and cascading diffusion decoders to generate high-resolution images based on the CLIP [17] text encoder. Imagen [26] focuses on language understanding by using a large T5 language model to better represent semantics. The latent diffusion model [22] improves computational efficiency by performing the diffusion process in low-dimension latent space with an autoencoder. Finally, Stable Diffusion (SD) [22] employs a cross-attention mechanism to inject textual conditions into the diffusion generation process, aligning with the provided textual input. However, it is difficult to provide precise action description in text, since user intent and machine understanding are not aligned. Furthermore, experimental results in Fig. 4 show that some actions are difficult to generate correctly without re-training, e.g., “performs a handstand”.
2.2 Controllable Action Generation
@@
In this paper, we focus on transferring the desired action to unseen people, characters, and even animals for photorealistic image generation. Existing efforts take source images and pose information (e.g., skeletal images or body parsing) as conditions to control the generation. Previous controllable solutions based on GANs [12, 13, 35] and VAEs [21, 32] suffer from training difficulties and poor generation results. Some subsequent works [24, 31] introduce text conditions to guide the action generation, yet fail with open vocabulary due to the small size of the vocabulary pools. Thanks to the significant advances of T2I diffusion models, recent methods [10, 14], in particular the popular ControlNet [34], involve additional trainable modules to add arbitrary conditions, improving the versatility and controllability. While gaining a tremendous amount of traction from the community, ControlNet refers to the provided skeleton image to generate the action, which reduces the flexibility and diversity. In addition, the objective of designing a general framework makes it not well-targeted to animals. In this work, we investigate customized solutions for action generation.
2.3 Subject-Driven Customization
@@
Due to the demand for generating images with user-specified subjects, customization methods [25, 5, 9, 29] tailored to the appearance have been studied in context of T2I generation. Specifically, DreamBooth [25] binds rare new words with specific subjects through fine-tuning the whole T2I generator. Textual Inversion [5] learns an extra identifier to represent the subject and adds the identifier as a new word to the dictionary of the text encoder. Custom Diffusion [9] only fine-tunes the key and value matrices of the cross-attention to represent new concepts. P+ [29] extends the textual-conditioning space with per-layer tokens to allow for greater disentangling and control. Despite the success achieved, the experimental results in Fig. 2 show their failure in action customization. A recent work ReVersion [7] makes progress in learning specific relations including some interactions from exemplar images. However, the design of the method, which specializes in learning spatial relations, makes it difficult to invert action information.
@@ 4.2 Action-Disentangled Identifier (ADI)
Expanding Semantic Inversion. To overcome the preference to low-level appearance features, we apply layer-wise identifier tokens to increase the accommodation of various features. Specifically, for the $l$-th layer where $l\in[1,L]$ and $L$ is the number of cross-attention layers in the T2I model, a new identifier token $v_{l}\in\mathbb{R}^{d}$ is initialized. Feeding the prompt with $v_{l}$ into the text encoder, the output tokens $\mathbf{y}_{l}$ can control the update of the latents in the $l$-th layer of SD, thus influencing the generation of the visual content. And the learned tokens from all layers can form a token set $\mathcal{V}$, which can then be paired with different subjects for generation. Rather than having a single identifier token take on the responsibility of reconstruction, having separate identifiers at different layers effectively ensures that more features are converted, including the action-related features we care about.
Learning Gradient Mask with Context-Different Tuple. The next step is to prevent the identifiers from learning features that are not relevant to the action and thus contaminating the subsequent image generation. Given $\mathbf{x}^{(a,c)}\in\mathcal{X}$ as an anchor sample, where $a$ denotes the specific action, and $c$ denotes the action-agnostic context contained in the image including human appearance and background, we can randomly sample another image $\mathbf{x}^{(a,\overline{c})}$ from $\mathcal{X}$, where $\overline{c}$ represents that the context is different from $c$. Taking the context-different tuple $\mathbf{x}^{(a,c)}$ and $\mathbf{x}^{(a,\overline{c})}$ as the input, we can calculate two gradients of the denoising loss $\mathcal{L}$ with respect to the identifier token $v$:
@@
where the semantic channels with a small difference can be regarded as action-related feature channels of the action $a$, which are expected to be preserved. Specifically, we sort the difference from the largest to the smallest, denoting the value at $\beta$ percent as $\gamma^{\beta}$ and taking it as a threshold. In other words, $\beta$ is the proportion of channels that are masked. Then, the mask that shares the same dimension as $v$ can be calculated. For the $k$-th channel,
@@
By overwriting the mask to the gradient of the anchor sample, the action-related knowledge is preserved and incorporated into the update of $v$, while the updates on action-agnostic channels are ignored. Note that since the specific visual invariance about the action changes slightly depending on the sample pair, the channels that are masked may not be exactly the same each time. Another point that worth highlighting is that both samples use the prompt of the anchor sample $\mathbf{x}^{(a,c)}$ when calculating the gradients. Since the visual context of $\mathbf{x}^{(a,\overline{c})}$ is inconsistent with the description in the prompt, the reconstruction loss favours larger gradients in the context-related channels. In this way, the action-related channels found through the threshold will be more accurate.
@@
Learning Gradient Mask with Action-Different Tuple. Although the context-different tuples have the same action semantics, there may be differences in the visualisation of the actions, and therefore the channels associated with the most representative action features do not necessarily have a smaller gradient difference. Since learning the gradient mask with the context-different only tuple is not stable and effective enough, we also construct action-different tuples to generate the gradient mask from another perspective. For each sample $\mathbf{x}^{(a,c)}$ in $\mathcal{X}$, we can use it to quickly train a subject-driven customization model (e.g., DreamBooth) that effectively inverts the most of the low-level context information. Therefore, by filling the prompt template of $\mathbf{x}^{(a,c)}$ with action descriptions that are different from $a$, the trained customization model can generate various action images as $\mathcal{X}^{(\overline{a},c)}$. Due to the one-shot training and the concise text, these images may not be consistent with the action descriptions, or the context may differ from the original $\mathbf{x}^{(a,c)}$, but in practice, we have found that the quality is sufficient to diversify the action variation. In this way, when $\mathbf{x}^{(a,c)}$ is sampled during training, we can randomly sample a image $\mathbf{x}^{(\overline{a},c)}$ from $\mathcal{X}^{(\overline{a},c)}$ to construct the action-different tuple. And the gradient of $\mathbf{x}^{(\overline{a},c)}$ with respect to the token $v$ can be calculated as
@@
where the semantic channels with small difference can be regarded as context-related feature channels of the action $a$, which are expected to be masked. Therefore, we have
@@
where $\lambda^{\beta}$ is the threshold here to mask $\beta$ percent of the channels.
Merging Gradient Masks for Context. According to the above discussion, the final input of the model is a triplet as =$\mathcal{I}=\left\{\mathbf{x}^{(a,\overline{c})},\mathbf{x}^{(a,c)},\mathbf{x}^{(\overline{a},c)}\right\}$. And two context masks can be created with different binary groups. Therefore, we can merge $m^{\overline{a}}$ and $m^{\overline{c}}$ to get the final context mask $m$. In practice, we keep only the intersection of the unmasked channels as unmasked, as we find this merging strategy performs better. Formally, we have
$m=m^{\overline{c}}\cap{m}^{\overline{a}}.$ (13)
Then, we overwrite $m$ to the gradient of the anchor sample:
@@ 5.1 Experiment Setup
Implementation Details. For our ADI, we set the masking ratio $\beta$ to 0.6 and use the AdamW [11] optimizer with a learning rate of 2e-4. For a fair comparison, we use 50 steps of the DDIM [27] sampler with a scale of 7.5 for all methods. Unless otherwise specified, Stable Diffusion v2-1-base is selected as the default pre-trained model, and images are generated at a resolution of 512$\times$512. All experiments are conducted on one A-100 GPU.
@@ 5.4 Ablation Study
We conduct ablation experiments on ActionBench to verify the individual effects of the proposed contributions. From the generation results in Fig. 5, it can be observed that (1) The removal of the extension to the semantic conditioning space diminishes the inversion ability of ADI. (2) Both the gradient masks learned from the context-different and the action-different tuples are essential. Removing either one can lead to inadequate learning of action knowledge or a degradation in the quality of the subject’s appearance. (3) We also attempt to reverse the gradient masks, i.e., updates to channels that should have been masked are preserved, and updates to other channels are cancelled. Obviously, this will result in action-related features not being inverted.
@@ 6 Conclusion
Acknowledgement
This work was supported by STI 2030—Major Projects (2022ZD0208800), NSFC General Program (Grant No. 62176215). This work was supported by Alibaba Group through Alibaba Research Intern Program.
@@ C.1 Comparison with Action-Prior DreamBooth
C.2 Additional Qualitative Results
To show the effectiveness of ADI, we illustrate additional generation results in Fig. 10, covering all actions within ActionBench.
Figure 10: Additional generation results by ADI.

### v3

@@ Abstract
This study focuses on a novel task in text-to-image (T2I) generation, namely action customization. The objective of this task is to learn the co-existing action from limited data and generalize it to unseen humans or even animals. Experimental results show that existing subject-driven customization methods fail to learn the representative characteristics of actions and struggle in decoupling actions from context features, including appearance. To overcome the preference for low-level features and the entanglement of high-level features, we propose an inversion-based method Action-Disentangled Identifier (ADI) to learn action-specific identifiers from the exemplar images. ADI first expands the semantic conditioning space by introducing layer-wise identifier tokens, thereby increasing the representational richness while distributing the inversion across different features. Then, to block the inversion of action-agnostic features, ADI extracts the gradient invariance from the constructed sample triples and masks the updates of irrelevant channels. To comprehensively evaluate the task, we present an ActionBench that includes a variety of actions, each accompanied by meticulously selected samples. Both quantitative and qualitative results show that our ADI outperforms existing baselines in action-customized T2I generation. Our project page is at https://adi-t2i.github.io/ADI.
@@
*Work done during internship at Alibaba Group.Corresponding author.
@@ 1 Introduction
To avoid the appearance leakage while accurately modeling the target action, we propose Action-Disentangled Identifier (ADI) to learn the optimal action-specific identifiers. Firstly, we expand the semantic conditioning space by applying layer-wise identifier tokens. Since existing works have analyzed that different layers have varying degrees of control over low-level and high-level features [30], such an expansion increases the accommodation of various features, making it easier to invert action-related features. Furthermore, we would like to decouple the action-agnostic features from the learning of action identifiers. To achieve this, we discover invariant mechanisms in the data that are difficult to vary across examples. Specifically, given an exemplar image with the specific action, another same-action image can be randomly sampled from the training data, forming a context-different pair. Meanwhile, leveraging mature subject-driven customization techniques, an image that shares the similar context can be quickly synthesized to form an action-different pair. To decouple the highly-coupled features, we disentangle action-agnostic features at the gradient level, and construct two context gradient masks by comparing the difference on the gradients over the input pairs. By overwriting the merged gradient mask to the gradient of the anchor image, the update of action-agnostic channels on the identifiers is discarded.
@@
  • We propose a novel action customization task, which requires learning the desired action from limited data for future generation. While existing customization focuses on reprinting appearances, we highlight this under-studied but important problem.
@@ 2 Related Work
Text-to-Image (T2I) Generation. Generating high-quality and diverse images from textual conditions has received considerable attention from both the research community and the general public. The previous dominant generative adversarial networks (GANs) [20, 31, 28, 38], consisting of a generator and a discriminator, suffer from unstable optimization and less diverse generations due to the adversarial training [2]. And variational autoencoders (VAEs) [8], which apply a probabilistic encoder-decoder architecture, are also prone to posterior collapse and over-smoothed generations [37]. Text-conditional auto-regressive models [18, 4, 3, 34] have shown more impressive results, but require time-consuming iterative processes to achieve high-quality image sampling. More recently, diffusion models have emerged as a promising alternative, achieving impressive results with open-vocabulary text descriptions through their natural fitting to inductive biases of image data [26, 19, 15, 22]. GLIDE [15] introduces text conditions into the diffusion process through the use of an unclassified guide. DALL-E 2 [19] employs a diffusion prior module and cascading diffusion decoders to generate high-resolution images based on the CLIP [17] text encoder. Imagen [26] focuses on language understanding by using a large T5 language model to better represent semantics. The latent diffusion model [22] improves computational efficiency by performing the diffusion process in low-dimension latent space with an autoencoder. Finally, Stable Diffusion (SD) [22] employs a cross-attention mechanism to inject textual conditions into the diffusion generation process, aligning with the provided textual input. However, it is difficult to provide precise action descriptions in text, since user intent and machine understanding are not aligned. Furthermore, experimental results in Fig. 4 show that some actions are difficult to generate correctly without re-training, e.g., “performs a handstand”.
@@
Controllable Action Generation. The paper focuses on transferring the desired action from examplar images to unseen people, characters, and even animals for photorealistic image generation. Existing efforts take source images and pose information (e.g., skeletal images or body parsing) as conditions to control the generation. Previous controllable solutions based on GANs [12, 13, 36] and VAEs [21, 33] suffer from training difficulties and poor generation results. Some subsequent works [24, 32] introduce text conditions to guide the action generation, yet fail with open vocabulary due to the small size of the vocabulary pools. Thanks to the significant advances of T2I diffusion models, recent methods [29, 10, 14], in particular the popular ControlNet [35], add arbitrary conditions to improve the versatility and controllability. While gaining a tremendous amount of traction from the community, ControlNet refers to the provided skeleton image to generate the action, which reduces flexibility and diversity. In addition, the objective of designing a general framework with additional trainable modules makes it not well-targeted to animals. In this work, we investigate customization solutions for action generation.
@@
Subject-Driven Customization. Due to the demand for generating images with user-specified subjects, customization methods [25, 5, 9, 30] tailored to the appearance have been studied in the context of T2I generation. Specifically, DreamBooth [25] binds rare new words with specific subjects through fine-tuning the whole T2I generator. Textual Inversion [5] learns an extra identifier to represent the subject and adds the identifier as a new word to the dictionary of the text encoder. Custom Diffusion [9] only fine-tunes the key and value matrices of the cross-attention to represent new concepts. P+ [30] extends the textual-conditioning space with per-layer tokens to allow for greater disentangling and control. Despite the success achieved, the experimental results in Fig. 2 show their failure in action customization. A recent work ReVersion [7] makes progress in learning specific relations including some interactions from exemplar images. However, the design of the method, which specializes in learning spatial relations, makes it difficult to invert action information.
@@ 4.2 Action-Disentangled Identifier (ADI)
Expanding Semantic Inversion. To overcome the preference to low-level appearance features, we apply layer-wise identifier tokens to increase the accommodation of various features. Specifically, for the $l$-th layer where $l\in[1,L]$ and $L$ is the number of cross-attention layers in the T2I model, a new identifier token $v_{l}\in\mathbb{R}^{d}$ is initialized. Feeding the prompt with $v_{l}$ into the text encoder, the output tokens $\mathbf{y}_{l}$ control the update of the latents in the $l$-th layer, thus influencing the generation of the visual content. And the learned tokens from all layers can form a token set $\mathcal{V}$, which can then be paired with different subjects for generation. Rather than having a single identifier token take on the responsibility of reconstruction, having separate identifiers at different layers effectively ensures that more features are converted, including the action-related features we care about.
Learning Gradient Mask with Context-Different Pair. The next step is to prevent the identifiers from inverting features that are not relevant to the action and thus contaminating the subsequent image generation. Given $\mathbf{x}^{(a,c)}\in\mathcal{X}$ as an anchor sample, where $a$ denotes the specific action, and $c$ denotes the action-agnostic context contained in the image including human appearance and background, we can randomly sample another image $\mathbf{x}^{(a,\overline{c})}$ from $\mathcal{X}$, where $\overline{c}$ represents that the context is different from $c$. Taking the context-different pair $\mathbf{x}^{(a,c)}$ and $\mathbf{x}^{(a,\overline{c})}$ as the input, we can calculate two gradients of the denoising loss $\mathcal{L}$ with respect to the identifier token $v$:
@@
where the semantic channels with a small difference can be regarded as action-related channels of the action $a$, which are expected to be preserved. Specifically, we sort the difference from the largest to the smallest, taking the value at $\beta$ percent $\gamma^{\beta}$ as a threshold. In other words, $\beta$of the channels are masked. Then, the mask that shares the same dimension as $v$ can be calculated. For the $k$-th channel,
@@
By overwriting the mask to the gradient of the anchor sample, the action-related knowledge is preserved and incorporated into the update of $v$, while the updates on action-agnostic channels are ignored. Note that since the specific visual invariance about the action changes slightly depending on the sample pair, the masked channels may not be exactly the same each time. Furthermore, both samples use the prompt of the anchor sample $\mathbf{x}^{(a,c)}$ when calculating the gradients. Since the visual context of $\mathbf{x}^{(a,\overline{c})}$ is inconsistent with the description in the prompt, the reconstruction loss favours larger gradients in the context-related channels. In this way, the action-related channels found through the threshold will be more accurate.
@@
Learning Gradient Mask with Action-Different Pair. Although the context-different pairs have the same action semantics, there may be differences in the visualization of the actions, and therefore the channels associated with the most representative action features do not necessarily have a smaller gradient difference. Since learning the gradient mask with the context-different only pair is not stable and effective enough, we also construct action-different pairs to generate the gradient mask from another perspective. For each sample $\mathbf{x}^{(a,c)}$ in $\mathcal{X}$, we can use it to quickly train a subject-driven customization model (e.g., DreamBooth) that effectively inverts the most of the low-level context information. Therefore, by filling the prompt template of $\mathbf{x}^{(a,c)}$ with action descriptions that are different from $a$, the trained customization model can generate various action images as $\mathcal{X}^{(\overline{a},c)}$. Note that this step is not necessary if the users can compile a dataset of varied actions by the same individual using pre-captured images. However, the data collection is usually arduous and lengthy, making fast training of a subject-driven customization a more convenient solution. Due to the one-shot training and the concise text, the generated images $\mathcal{X}^{(\overline{a},c)}$ may not be consistent with the action descriptions, or the context may differ from the original $\mathbf{x}^{(a,c)}$, but in practice, we have found that the quality is sufficient to diversify the action variation. In this way, when $\mathbf{x}^{(a,c)}$ is sampled during training, we can randomly sample a image $\mathbf{x}^{(\overline{a},c)}$ from $\mathcal{X}^{(\overline{a},c)}$ to construct the action-different pair. And the gradient of $\mathbf{x}^{(\overline{a},c)}$ with respect to the token $v$ can be calculated as
@@
where the semantic channels with small difference can be regarded as context-related channels of the action $a$, which are expected to be masked. Therefore, we have
@@
where $\lambda^{\beta}$ is the threshold here to mask $\beta$of the channels.
Merging Gradient Masks for Context. Due to the noise introduced by context variations, identifying action-relevant channels using only context-different or action-different pairs would be difficult and unreliable. As an evidence, the average overlap rate of channels preserved by both masks at each training step is 30.26%. Therefore, given the input triple =$\mathcal{I}=\left\{\mathbf{x}^{(a,\overline{c})},\mathbf{x}^{(a,c)},\mathbf{x}^{(\overline{a},c)}\right\}$, we can merge the two obtained masks $m^{\overline{a}}$ and $m^{\overline{c}}$ to get the final context mask $m$. In practice, we keep only the intersection of the unmasked channels as unmasked, as we find this merging strategy performs better. Formally, we have
$m=m^{\overline{c}}\cap{m}^{\overline{a}},$ (13)
which is overwritten to the gradient of the anchor sample:
@@ 5.1 Experiment Setup
Implementation Details. For ADI, we set the masking ratio $\beta$ to 0.6 and use the AdamW [11] optimizer with a learning rate of 2e-4, while the training takes 3000 steps. For efficiency, DreamBooth for action-different pairs does not generate class-preservation images. While only one image is used for training, the initial learning rate is 1e-6, and the training takes 2000 steps. For a fair comparison, we use 50 steps of the DDIM [27] sampler with a scale of 7.5 for all methods. Unless otherwise specified, Stable Diffusion v2-1-base is selected as the default pre-trained model, and images are generated at a resolution of 512$\times$512. All experiments are conducted on a NVIDIA A100 GPU.
5.2 Quantitative Comparison
@@ 5.4 Ablation Study
We conduct ablation experiments on ActionBench to verify the individual effects of the proposed contributions. From the generation results in Fig. 5, it can be observed that (1) The removal of the extension to the semantic conditioning space diminishes the inversion ability of ADI. (2) Both the gradient masks learned from the context-different and the action-different pairs are essential. Removing either one can lead to inadequate learning of action knowledge or a degradation in the quality of the subject’s appearance. We attribute this to the fact that learning from a single pair is inherently noisy due to varied action visuals and the interference of action-irrelevant information. (3) We also attempt to reverse the gradient masks, i.e., updates to channels that should have been masked are preserved, and updates to other channels are cancelled. Obviously, this will result in action-related features not being inverted.
@@ 6 Conclusion
Acknowledgement This work was supported by STI 2030—Major Projects (2022ZD0208800), NSFC General Program (Grant No. 62176215). This work was supported by Alibaba Group through Alibaba Research Intern Program.
@@ Appendix B Baseline Details
All baselines use the prompt template provided by the ActionBench. Each prompt details its image content, leaving the action blank for filling with identifiers from different methods. Other details are:
@@ C.1 Comparison with Action-Prior DreamBooth
C.2 Generalization Across Diverse Styles
ADI is designed to separate and inverse abstract the action concepts from the details of subjects and objects, background, color, or style in user images. This allows the generation images to generalize to specific styles through prompting, shown as Fig. 10.
[…]
C.3 Visualization of Cross-Attention Maps
To explain why certain channels can be interpreted as “action-related”, we visualize the cross-attention maps related to the learned identifiers in Fig. 11. As observed, the learned identifiers focus more on the contour information of the actions rather than the human body. This indicates that ADI avoids reversion on appearance information, thereby enabling generalization to different subjects.
Figure 11: Visualization of cross-attention maps associated with the learned identifiers.
C.4 Visualization of Action-Different Pairs
We present the generated images in the action-different pairs in Fig. 12 for reference. Using only a single image for training, the subject-driven model can change actions while preserving contextual information as much as possible. Although the quality of the image may be insufficient, it does not hinder the final inversion of action knowledge.
Figure 12: Visualization of subject-driven generation results for action-different pairs.
C.5 Additional Qualitative Results
To show the effectiveness of ADI, we illustrate additional generation results in Fig. 13, covering all actions within ActionBench. The generated images maintain the same action while offering a rich diversity, indicating that the learned identifiers contain solely action information and do not encapsulate irrelevant contextual details such as background, appearance, or even orientation.
Figure 13: Additional generation results by ADI, encompassing all the actions within ActionBench.